B1 Run Time = 80 sec/unit Station C Station ACo- /O = 170 sec Run time = 35 sec/unit CO = 72 sec Run Time 38 sec/unit CO 80 sec Station B2 Run Time = 80 sec/unit C/O 170 sec For the process shown above, calculate the raw processing times (in seconds) for the following batch sizes: N=2 N-6 N 12 Raw processing times (enter whole numbers)
Answer #1

For N = 2

Total processing time = Run time*2 + Change over time for each station

Station A = 2*35 + 72 = 142

Station B1= 2/2*80 + 170 = 250 (Half the batch size units considered)

Station B2 = 2/2*80 + 170 = 250 (Half the batch size units considered)

Station C = 2*38 + 80 = 156

Total = 142+250+250+156 = 798 sec

For N = 6

Total processing time = Run time*2 + Change over time for each station

Station A = 6*35 + 72 = 282

Station B1= 6/2*80 + 170 = 410 (Half the batch size units considered)

Station B2 = 6/2*80 + 170 = 410 (Half the batch size units considered)

Station C = 6*38 + 80 = 308

Total = 1410 sec

For N = 12

Total processing time = Run time*2 + Change over time for each station

Station A = 12*35 + 72 = 492

Station B1= 12/2*80 + 170 = 650 (Half the batch size units considered)

Station B2 = 12/2*80 + 170 = 650 (Half the batch size units considered)

Station C = 12*38 + 80 = 536

Total = 142+250+250+156 = 2328 sec